Find two numbers whose difference is 4 and the difference of their squares is 88.
Answer by lwsshak3(11628)

Find two numbers whose difference is 4 and the difference of their squares is 88.
**
let x=one of 2 numbers
let y=other number
..
x-y=4
x=y+4
..
x^2-y^2=88
(y+4)^2-y^2=88
y^2+8y+16-y^2=88
8y=88-16=72
y=72/8=9
x=y+4=13
numbers are: 9 and 13